OpenStudy (anonymous):

what is the boundary of {1/n:n in N}

OpenStudy (anonymous):

is it (0,1)??

OpenStudy (anonymous):

Hmm, what is the definition of the boundary in this case?

OpenStudy (anonymous):

well I thought boundary is just a point which has a neighborhood that contains points in the set and not in the set.

OpenStudy (anonymous):

I would think the boundary would be maybe \[ \{0\}\cup \{1\} \]

OpenStudy (anonymous):

why would it be that and not include 1/2, 1/3, 1/4... etc.?

OpenStudy (anonymous):

Oh, I see what you mean...

OpenStudy (anonymous):

The boundary can't be in the interior, correct?

OpenStudy (anonymous):

correct. this set does not have any interior points because every neighborhood of x must contain every point in the set.

OpenStudy (anonymous):

Then the boundary is just the set itself.

OpenStudy (anonymous):

so I would just say bd = {1/n}

OpenStudy (anonymous):

\[ \partial \left\{\frac 1n:n \in N\right\} = \left\{\frac 1n:n \in N\right\} \]